Brahms -- September activity report

From: Flemming Videbaek (videbaek@sgs1.hirg.bnl.gov)
Date: Wed Sep 18 2002 - 12:30:29 EDT

  • Next message: Flemming Videbaek: "Brahms minimeeting"

    Activity Report September  16 2002.
    --------------------------------------------
    
    Dear Collaborator,
    Included are the activity progress report from each institution -
    The contributions are essentially posted as received -
    The plan is to have the next report in about 3-4 weeks times;
    If you want more details on specific subjects please feel free to contact
    the relevant persons
    (initial of correspondent is given after each contribution).
    
    best regards
        Flemming
    
    
    TAMU BRAHMS progress report 9-Sep-2002
    
    Analysis on pp data continues. We are concentrating on extracting
    negative/positive particle ratios. To date a "reasonable" sample of data
    at all angle and field settings has been reduced into "DST" form.
    
    The following is the analysis chain in use right now:
    1. Local track reconstruction on the CRS farm
    2. Matching on CAS farm
    3. DST on CAS farm
    4. Generate 3D spectrum y-pt-mass2 with conditions of interest
    5. Analyze projections of that spectrum.
    
    The following are angle/field settings that are currently being used.
    Arm angle current (amps D1; FS, D5; MRS) field (fraction FS, kG MRS)
    FS 3 1124 1/3 field
    FS 4 1124 1/3 field
    FS 12 843 1/4 field
    MRS 35 700 4 kG
    MRS 45 700 4 kG
    MRS 60 500 3 kG
    MRS 90 1000 6 kG
    
    These runs were picked because they are the only ones that have both
    polarities of the same angle and field setting. In addition to the
    above, there are a number of MRS run pairs where the angle and field
    were identical (field had opposite polarities), but in one run the MRS
    was at the nominal position and in the other the MRS was positioned 50
    cm back. This data can presumably be used because the acceptance
    difference between those two settings should be well known. In addition,
    there are some runs where we have pairs for both the nominal and 50 cm
    positions of the MRS. These have not been included yet.
    
    Shown on
    http://Cyclotron.tamu.edu/hagel/brahms/ProgressReport-09-Sep-2002/ppYPt.gif
    is the y-pt coverage achieved with this set of runs.
    
    Ratios were then extracted as a function of pt for the different
    settings. These ratios are shown at:
    3 deg
    http://Cyclotron.tamu.edu/hagel/brahms/ProgressReport-09-Sep-2002/spectra3De
    gFS1124PtRatio.gif
    4 deg
    http://Cyclotron.tamu.edu/hagel/brahms/ProgressReport-09-Sep-2002/spectra4De
    gFS1124PtRatio.gif
    12 deg
    http://Cyclotron.tamu.edu/hagel/brahms/ProgressReport-09-Sep-2002/spectra12D
    egFS843PtRatio.gif
    35 deg
    http://Cyclotron.tamu.edu/hagel/brahms/ProgressReport-09-Sep-2002/spectra35D
    egMRS700PtRatio.gif
    45 deg
    http://Cyclotron.tamu.edu/hagel/brahms/ProgressReport-09-Sep-2002/spectra45D
    egMRS700PtRatio.gif
    60 deg
    http://Cyclotron.tamu.edu/hagel/brahms/ProgressReport-09-Sep-2002/spectra60D
    egMRS500PtRatio.gif
    90 deg
    http://Cyclotron.tamu.edu/hagel/brahms/ProgressReport-09-Sep-2002/spectra90D
    egMRS1000PtRatio.gif
    
    Several brief comments can be made about these plots. The first is that
    there has been a persistent discrepancy between ratios generated from
    the FFS and ratios generated from the BFS which has not been fully
    understood. Looking at these ratios as a function of pt for the FFS and
    BFS was driven by this discrepancy to see if the differences could be
    explained by a dependence of the ratios on pt. We see at the 3, 4 and 12
    deg settings that a number of the differences are explained to within
    error bars by looking at overlapping pt bins. Nevertheless, some of the
    ratios continue to be different. There was some possibility that these
    were introduced by a bug in the generation of DSTs. The bug was
    corrected and dsts generated again during the initial writing of this
    document. Very little change was found. The bug had to do with the FFS
    track id being returned when asking for the BFS track id. This could be
    important in the Au + Au run, but as there is very seldom more than one
    track in pp, most of the track ids are the same, namely 0.
    
    Another comment concerns the 35 deg setting. These ratios are
    systematically lower than the other ratios and do not follow any kind of
    trend as will be shown below. The particular runs for that angle/field
    setting (namely runs 6270-631) came during a time where the number of
    trig 3 events in coincidence with the inelastic triggers was very
    depressed even though there are trig 3 events as well as INEL triggers.
    This is not yet understood, but to get around the problem, the vertex
    for those runs was determined by projecting the MRS track back to the
    beamline. In runs where both were working, a correlation is seen. In
    fact, it is the same technique that is used to calibrate the INEL
    vertex. This was done and ratios are more in line with trends, but it
    seems not fully. This is still under study. (added note /fv- a timing shift
    occurred
    For these runs causing tdc signals to shift outside a cut window)
    
    If the pt spectra are integrated from 1 GeV/c to 4 GeV/c and the ratios
    plotted as a function of rapidity, we obtain the plot shown at:
    http://Cyclotron.tamu.edu/hagel/brahms/ProgressReport-09-Sep-2002/ratiosVsY.
    gif
    
    It is seen in this plot that the third point from the left, the 35 deg
    MRS setting, is depressed relative to the trends seen. As pointed out
    above, this is still under study. Aside from that point, the trends seem
    to show some smooth trends. Assuming no flaws are found in the analysis,
    in particular particle id, these ratios are the current "best" and could
    presumably soon be used to start to interpret the data.
    
    This is an overview of what is coming out of the pp analysis at TAMU to
    date. This report skips a lot of minute detail. The other "important"
    project is that a report is being prepared which details what has been
    done as well as steps taken to achieve quality control. This report will
    be distributed as soon as it is far enough along. In particular it is
    meant to "convince" the reader that the above presented data is real.
    
    It should be noted that all plots shown here are to be treated as
    preliminary and are not to be distributed until the collaboration has
    had a chance to review and approve them.
    /kh
    
    Kraków group progres raport.
    
    Construction of the HV splitters has been completed. T. Kozik and P. Staszel
    will arrive to the BNL on Sept. 23 to install these new devices and to
    prepare the DC's for coming run.
    
    Z. Majka gave a talk on the BRAHMS results at the Light Cone 2002
    Meeting at Los Alamos.
    
    /zm
    
    
    KANSAS
    
    The design and construction of the C4 detector is proceeding. The final
    design for the
    detector box and mirrors can be found at
    http://www.phsx.ukans.edu/~sanders/Cherenkov.
    The front box has been constructed and is being tested for leaks. The back
    box containing the
    large 45 deg. mirrors has been contracted out to a local shop. This box has
    a promised completion
    date of Oct. 11. A prototype has been constructed for the PMT + pyramid
    mirror mount assembly.
    In addition, all of the PMT mounting hardware has been completed. This
    prototype will be
    shipped to BNL within the next few days for testing with the actual PMT
    tubes. Construction of
    the actual PMT+pyramid mirror mounts is expected to start next week in the
    Kansas shops and is expected to be completed in about two weeks. BNL
    (Ramiro) is working on acquisition of the
    glass pieces (mirrors and windows) needed for the detector. The detector
    mount will also be
    developed at BNL.
    
    Hiro successfully defended his thesis on the Brahms charged-particle
    multiplicity measurement.
    He is still busy with multiplicity analysis, however, as we start to look at
    the pp data obtained with the
    Si+Tile arrays. One question is whether we will want to keep the Si array in
    place for the light-ion
    runs.
    
    /sjs
    
    
    
    BNL
    
    
    The shielding for the FS is now complete. The RICH electronics have been
    moved to the rack on the opposite side of the spectrometer. The H1
    electronics have been removed in preparation of relocating.
    
    The prototype trigger board has arrived at BNL.  The main purpose of the
    unit is to form a dual coincidence i.e. for up/down tubes on slats. A module
    can handle 16 slats. The modules put out and OR as well as daisy chain for a
    "Grand OR". Individual slats can be "turned off" by a bit mask. Also entire
    boards can be vetoed by an input. It has been tested and there are minor
    updates for the final version.  The boards will be used to form triggers for
    TOFW, spectrometer trigger counters and most likely H1 and H2. It is
    expected that the updated version will be sent to the vendor to fabricate
    for the experiment within one week. The insertion time to convert the signal
    to ECL and to form a "grand OR" is 43 ns. A Camac crate is required for the
    NIM to ECL conversion modules (existing units) and a VME crate for the new
    coincidence modules. It is planned that one pair of crates will be on the
    MRS for TOFW and one in the FEH for H1, H2, and spectrometer counters.
    
    /db
    
    Work on TPCs:
    The Mylar windows installed in all TPCs turned out to be the biggest source
    of contamination of the gas. The drift velocity monitors saw a difference in
    drift velocity upstream and downstream of the TPC, that difference was as
    high as 20% in TPM2. The changes in drift velocity had some correlation with
    atmospheric pressure but not a one to one. The change in drift velocity
    measured with the monitors correlated very well to the vertex variations as
    calculated with tracks from TPM1 (and TPM2).
    A 1 mil thick Mylar foil has a measured water vapor transmission rate of : 2
    g/(100 square inches 24 hours) and permeability to oxigen molecules equal to
    6 cc/(100 square inches 24 hours) at 25 degrees centigrade and 1 atm.
    Kapton would let more water vapor in (3.5 g/100 square in. 24 hr ) but
    aluminum foils are a much better solution, the only concern was the
    establishment of a ground plane close to the drift cathode. I have decided
    to change all Mylar foils to 1 mil aluminum. These foils are all grounded.
    With aluminum foils the measured values of drift velocity are 1.7
    cm/microsec upstream and 1.67 cm/microsec downstream
    Both measurements correlate well with barometric pressure.
    T1 has had high voltage on the drift cathode for almost one week with no
    breakdown, the current drawn from the power supply is up by 0.08 % I
    consider that negligible. The electric field inside the field cage should
    only see small changes that extend into the cage by roughly one strip width
    and the field lines that left the volume and went to infinity end up in the
    new well defined ground plane of the Al foils.
    
    TPM2 and T1 are ready to take data. TPM1 still has aluminized Mylar foils
    that will be changed to pure Al, T2 will also have its foils changed.
    All broken wires in T1 and T2A have been replaced. One wire "broke" in T2B
    when the wire plane was reinstalled on the TPC.
    While that wire was being removed, three more wires "broke". After
    inspecting the wires I think that they do not break, they come loose under
    the solder. The wire frames are not rigid enough, and every time we handle
    them we may loose more wires.
    I decided to put the frame back on the TPC and not to fix the missing wires.
    This is the location of the wires:
    Two anodes over pad row 8
    One over pad row 6
    and one over pad row 3
    (Pad row one is closer to the vertex)
    
    The drift velocity monitors will be read with Camac TDCs to avoid by losses
    of data in the Fastbus TDC (unexplained and intermittent effect)
    /RD
    
    Over the summer all the pii's have been upgraded to RH7.3. All code in the
    operator account is modified to work with gcc 3.04. Various upgrades to
    firewall, access controls imposed by lab security has been done. This has
    not all gone smoothly, problems remains with ssh on opus, and nfs stale file
    handles seems to appear more often than before. The opus data disk were
    upgraded and the spool disk is now 400Gb. During an RCF HPSS stress testing
    a few weeks back this new disk storage worked without any problems.
    
    Readout of the DVM time-signals has been moved from a FB pipeline TDC to a
    Camac to see if this fixes an intermittent readout problem.
    
    The performance of timing counter for FS (TD1) to be used in dA and pp was
    re-examined based on the pp runs in January. A brief report can be found on
    the pp counter page.
    http://www.sdcc.bnl.gov/brahms/private/detectors/pp/index.html
    
    /fv
    
     The Lambda feed-down correction has been estimated for protons  in MRS and
    FS.
     About 75% of protons from Lambdas are estimated to be recognized
     as protons in MRS. These calculation was made with HIJING data.
     Cut values used for selecting protons (+-3sigma in z and y) are
     from the actual data.  No significant vertex and angle dependence
     observed. For FS, about 80% of protons from Lambdas survive the cuts.
     More detailed information will be posted by Eun-Joo in a few days.
    
    Calibrations has been done for the earlier  run 4532 -  run 5357 , namely:
    BB counter Adc Gain, Delta Tdc, Slewing, VtxOffset were calibrated for most
    runs. Values are stored inh DB.
    For these runs one pass of the MRS reduction (for PID) was done, following
    the earlier tracking, though i t
    have to done again. This in part since tofw and zdc calibrations are needed
    for all these runs.
    ejk/
    
     High-pt suppression has been looked at for 40 degree data.
     By looking at the ratio of (0-10%)/(40-50%), we see very similar
     trend (turn over at ~ pt = 2GeV/c) as observed at y=0.
     (See the figure
    
    http://pii3.brahms.bnl.gov/~jhlee/brahms/analysis/plots/40degree_high_pt.ps)
     The ratios of longitudinal momentum pl also plotted in the figure
     showing similar trend.  There seems to be a possibly interesting
     observation in the data. The ratios of pt/pl for peripheral (40-50%)
     is higher then central (0-10%) at p<2GeV/c.
     (See the figures
     http://pii3.brahms.bnl.gov/~jhlee/brahms/analysis/plots/40pt_pl_h_minus.ps)
     http://pii3.brahms.bnl.gov/~jhlee/brahms/analysis/plots/40pt_pl_h_plus.ps)
     This might be a reflection of centrality dependent radial flow/slope
    changes or centrality dependent reconstruction efficiencies.
     Statistics is rather poor since it's from one field
     setting (700A) not to introduce any possible bias from mixing different
     settings. We have about factor of 4-5 more statistics at 40 degree.
    
     HIJING data with baryon junction (HIJING/B) were looked at to see
     how the code describes our data.  (See the figure
    
    http://pii3.brahms.bnl.gov/~jhlee/brahms/analysis/net_proton_with_models.ps)
     The data is for top 10% and AMPT is for 10% centrality, HIJING and HIJING/B
     is for impact parameter 0-4.5fm)
     Around mid rapidity (y = 0 - 2), HIJING/B gives more stopping as expected
     and overlaps with AMPT but it gives lower values at higher rapidities.
    
    /jhl
    
    Oslo/Bergen
    
    Spectra and yields from Au-Au collisions - Jens Ivar:
    
    Currently working on:
    * Putting together a presentation for the PANIC conference,
    titled "Rapidity dependence of particle ratios and spectra at
    $\sqrt{s_{NN}}$ = 200 GeV Au-Au collisions" and focussing on
    strange particle production.
    * Gaining experience with BRAG - learning how to include new
    detector elements, running BRAG and subsequent processing of
    simulated data.
    
    Future plans:
    * Continue working on hadron spectra and K/pi ratios for the
    higher rapidities
    * Recheck various corrections
    
    _______________________________________________________________
    
    Pi0 measurements in BRAHMS - Zhong-Bao Yin
    
    Currently working on:
    * Gaining experience with BRAG - learning how to include new
    detector elements, running BRAG and subsequent processing of
    simulated data.
    
    Future plans:
    * Incorporate a photon converter (lead block in front of TPM1)
    in Brag and perform simulation studies
    
    _______________________________________________________________
    
    Efficiency calculations - Truls:
    
    * Truls has rewritten the efficiency code BEAT to a bratmain compatible
    framework, using "package" classes BrEffGeantModule and BrEffRecoModule
    to take care of the multiple inputs and track tables.  The programming
    problem has finally been solved!  :-)
    
    Currently working on:
    * Produce tracking efficiency results for MRS and FFS!
    
    Future plans:
    * Study the TPC two-track resolution
    
    ___________________________________________________________
    
    Lambda project - Elin, Bjørn, Trine:
    
    * Several bugs in the DST to V0 candidate code identified between
    last progress report and QM.  Another bug identified after QM, and
    now we have got rid of the ugly peak at ~1.35 GeV in the invariant
    mass spectrum, and a much clearer lambda signal.
    See:
    http://lynx.uio.no/trine/brahms/lambda/lambda_35_decpt1.ps
    http://lynx.uio.no/trine/brahms/lambda/lambda_35_decpt2.ps
    for invariant mass spectra (35 deg) before and after the bug was
    corrected.
    
    Currently working on:
    * Lambda search at more angles, utilizing as much as possible of
    the available data set
    * Incorporating acceptance code for lambdas in Peter's acceptance
    framework.
    
    Future plans:
    * Extract Lambda and anti-Lambda ratios and yields for the MRS
    at selected angles.
    
    ___________________________________________________________
    
    Data from pp run - Mads, Bjørn:
    
    Currently working on:
    * Gaining experience with the bdst analysis chain for MRS and FS
    
    Future plans:
    * Extract K/pi yields (Mads) and net protons (Bjørn)
    
    /tt
    
    
    
    NBI
    
    Here at NBI we are presently working on several important and exciting
    projects, what follows is a brief account of several of them.
    1. FS efficiency.  Pawel is leading the effort to understand the
    efficiency of the FS.  There are still a few outstanding issues, but
    progress is being made.  There will be a small meeting on this topic
    at BNL on 2. Oct., which will certainly lead to a resolution of all
    open problems.
    2. Mads Gammeltoft completed his Master's.  Congratulations to Mads!
    3. Christian is busy writing his Master's thesis.  So do not bother
    him with other things (unless they are really, really important).
    4. Quark Matter proceedings.  Djamel is nearly finished with his
    contribution. Claus has finished his, while Ian is lagging a bit
    behind, but hopes to have a draft very soon.  I have received several
    requests for our data, so it seems that our QM show was well received.
    Congratulations to all of us!
    5. Peter and Djamel are looking into the y=3 data, to try to
    understand some things which were not fully understood at the time of
    QM. One result of this work is that a bug in the FFS-BFS matching(which was
    present in the
    code used to obtain the QM results)  was found. Note that this is related to
    point 1
    above.   One of the results will be a short write-up of the methods.
    5a) We will (quickly!!) also redo the other analyses which were done
    for QM as well as extend these to other pt,y regions for which we have
    data. It is of particular interest, at least to us, to have better
    rapidity coverage for the net--proton rapidity density, not only at
    the highest available rapidities, but also in the gaps between the
    data presented at QM.
    6. Peter is preparing his talk for PANIC, and a first draft will be
    circulated to the collaboration for comments early next week.
    7. Claus has been playing with Hijing for comparisons with the results
    coming from point 5.  We also have plans to look at d+Au simulations
    to get some idea what to expect.  Has anyone else looked at this, and
    if so what good advice do you have?
    8. Claus is also revisiting some of the Cherenkov analysis code, with
    hopes of improving the obtained results.
    9.  We are beginning to discuss how to improve the global tracking in
    the FS.  It is too early to say more, but anyone who has ideas should
    contact me, or write directly to the "dev" list.
    
    In addition to this we have a visitor from Giessen, Jim Ritman, who is
    looking at the p+p data.  He managed to reconstruct lambdas from
    events in which both a pion and proton went through the full MRS.
    Nice work, Jim!
    Finally, if anyone is interested in looking at data but needs help
    getting started, please just ask on the list and one of us will be
    happy to (at least try to) help you get started.  We have a wealth of
    beautiful data, let us analyze (and write it up!!) it in a timely
    manner.
    /igb
    



    This archive was generated by hypermail 2b30 : Wed Sep 18 2002 - 12:21:45 EDT